OpenStudy (anonymous):

HELP! HELP! HELP! HELP! HELP! I CAN'T UNDERSTAND THIS ONE CONCEPT,AND IF I DON'T, I'm PROBABLY GOING TO FAIL MY ONLINE QUIZ! I CAN'T FIGURE OUT HOW TO USE EVERYTHING ON THE EQUATION EDITOR TO EXPLAIN MY PROBLEMS...AND I CAN'T FIGURE OUT HOW TO TYPE MY PROBLEMS FROM THE KEYBOARD! MY TEXTBOOK AND SOLUTIONS MANUEL DON'T MAKE ANY SENSE ON THE CONCEPT, TO ME ANYWAY! MY MOM CAN'T HELP ME...HEEEEELLLLLLLP!!!!!!!!

hero (hero):

Is this real?

OpenStudy (amistre64):

that question aint gonna help either, cant make hide nor hair of what it is you need

hero (hero):

Just post the question

OpenStudy (anonymous):

Why does t + 5^2 squared = t + 5, and a^2 + 6a + 9 squared =|a + 3|? I can't find any pattern for concerning when to use absolute symbols for square root answers...please help!

OpenStudy (anonymous):

You know those moments when math brought you to tears? I'm having one of those moments right now, pretty much..

hero (hero):

You have to find a pattern for this? What was the original question?

hero (hero):

No, sorry, math never brought me to tears...at least not yet

OpenStudy (anonymous):

I figure out how to do math problems by trial and error, generally...unless somebody can explain it good to me....I call my method of doing math 'finding patterns'...instructions for both problems I tried my best to post...'Find each square root. Assume that all variables are unrestricted, and use absolute value symbols when necessary.

OpenStudy (anonymous):

then It just gave me the (t + 5)^2 squared and a^2 + 6a + 9 squared as a couple of the many problems in the section....

hero (hero):

Hmmm

hero (hero):

What is your email address? I'm going to post a classroom

OpenStudy (anonymous):

indianaelena@gmail.com

OpenStudy (anonymous):

where will you be posting this classroom?

hero (hero):

in your email

OpenStudy (anonymous):

not seeing yet.

OpenStudy (anonymous):

still not there.

OpenStudy (anonymous):

it's in now.

OpenStudy (anonymous):

hero..don't have a microphone...for my computer..